# Multimodal Conditional 3D Face Geometry Generation

### v1

@@ 4. Results
We now show several results and applications of our new multimodal conditional 3D face geometry generation method. We begin by demonstrating control over the generated facial geometry using FLAME’s identity and expression parameters (Section 4.1). We then demonstrate multimodal conditioned geometry generation by guiding the denoising process using sketches, sparse 2D landmarks, Canny edges and text. We evaluate the effectiveness of these different modalities in guiding the generated geometry both qualitatively and by using quantitative metrics in Section 4.2. In addition to using different conditioning modes, we show how one can also spatially restrict the guidance to a particular face region to perform precise geometry edits in Section 4.3. Finally we generate dynamic facial performances by guiding our model from video inputs and demonstrate that our method can produce facial animations that are stable across time (Section 4.4).
4.1. Identity and Expression Parameter Conditioning
@@
The base conditioning used to train our geometry generator are the identity and expression parameters from the FLAME model (Li et al., 2017). We use 300 identity parameters (), 100 expression parameters () and 3 jaw pose () parameters. We combine these FLAME parameters into a 403-dimensional conditioning vector which we pass through a 3-layer MLP with Leaky ReLU activation functions prior to injecting it into the diffusion model via cross-attention.
@@
Recollect that we do not use the geometry from the FLAME model itself to train our diffusion model. Instead we fit the FLAME model to the high quality facial geometry from Chandran et al. (2020) only to obtain identity and expression parameters, and train the diffusion model directly on the geometry captured by Chandran et al (2020). As there is a loss in geometric detail during this FLAME fitting step, our diffusion model is expected to compensate for the information missing in the FLAME parameters, but present in the ground truth facial geometry. In other words, our geometry generator has to re-interpret the FLAME parameters such that it can re-produce the training data and therefore does not memorize the FLAME parameter space exactly.
@@
Disentangling Identity and Expression. Our diffusion model also preserves the disentanglement between facial identity and expression that is present in FLAME. In Fig. 5, we show how a smooth interpolation of FLAME’s identity and expression coefficients results in a smooth, yet nonlinear, interpolation of our generated geometry. We can observe that the facial expression remains fixed when interpolating between identities, and vice versa (please also refer to the supplemental video). To eliminate the randomness in the generation, we used the same initial noise to generate the interpolated geometries along with DDIM sampling (Song et al., 2021).
Extreme Expressions Due to the limited number of extreme expressions in our training datasets, our model can produce suboptimal results when conditioned on FLAME parameters corresponding to extreme expressions (Wide open mouth for example). However, this is purely a data limitation and could be resolved by having a larger and more balanced dataset, by oversampling expressions during training or by weighting the loss towards focusing more on extreme expressions.
4.2. Multimodal Conditioning
Beyond the underlying FLAME-based control, we introduce additional conditioning modes to control our diffusion model following Section 3.4. We now discuss the results of facial geometry generation by conditioning our diffusion model on sketches, sparse 2D landmarks, Canny edges, portrait RGB photos and text. To create facial geometry from text input, we leverage the fact that CLIP (Radford et al., 2021) embeds both RGB images and text into a shared latent space. Therefore, instead of training separate cross-attention layers for text, we re-use those learned from our portrait images and directly feed them with the CLIP vectors obtained from a text prompt.
Geometries generated for different conditionings from each of these additional modes are shown in Fig. 6. While our generated geometry follows the identity and expression seen in the input conditioning signal, the degree to which the geometry can match the conditioning varies from mode to mode. For example, as portrait photos contain the most amount of information about a face when compared to sketches, Canny edges, landmarks or text, geometries generated by conditioning our model on portrait images tend to more tightly constrain the generated identity and expression (also see Table 2).
@@ 4.2. Multimodal Conditioning
| No augmentations | 4.093 | 3.692 | 2.170 |
| With augmentations | 3.757 | 3.352 | 2.085 |
| V2V error | Mean | Median | Std |
| 2D landmarks | 6.489 | 5.765 | 2.314 |
| Canny edges | 5.938 | 5.488 | 1.897 |
| Sketch | 5.630 | 5.034 | 2.099 |
| Portrait photo | 5.394 | 4.874 | 2.025 |
| FLAME parameters | 4.772 | 4.462 | 1.317 |
@@
To control the strength of the guiding condition, we make use of classifier-free guidance (Ho and Salimans, 2021) following Eq. 6. Fig. 7 demonstrates the effect of different guidance strengths for each condition. Increasing the guidance strength increases the effect that the conditioning has on the resulting geometry. For example, in Fig. 7, the expression of the generated geometry of the subject in the first row displays stronger wrinkles, and a closer match to the portrait image when setting compared to setting it to . Finally, despite training on purely studio data, we show how our model responds to conditionings derived from in-the-wild data in Fig. 8.
@@
4.2.1. Quantitative Evaluation
To evaluate the effectiveness of each of our conditioning modes in guiding the generated geometry towards ground truth scans, we compute the Euclidean error between the generated geometry and the ground truth geometry for each conditioning mode. In Table 2, we report the vertex-to-vertex (V2V) error of 178 shapes from our validation set for each type of conditioning. We observe that conditioning signals that are more descriptive, such as FLAME parameters or portrait photos, obtain a lower error when compared to signals that are less descriptive of the final geometry (2D landmarks, Canny edges and sketches). For the base FLAME parameter conditioning, we also visualize the error maps on the geometry in Fig. 9. We also highlight the importance of training with 3D geometric augmentations in Table 1 where our diffusion model trained with augmentations outperforms one trained without data augmentations, when evaluated on only neutral expressions from our validation set.
Finally, as our method involves learning additional modes of conditioning on top of an underlying FLAME conditioned diffusion model, we can also use more than one conditioning signal at inference time to guide the generation according to Eq. 6. In Fig. 10, we show how combining both FLAME parameter and portrait image conditioning lowers the vertex error on a validation sample, as the denoising UNet now has access to more information about the desired identity and expression.
4.3. Geometry Editing
The latent space of our autoencoder preserves the spatial layout of the original UV position map, much like how the latent space of the image autoencoder in text-to-image models (Rombach et al., 2022) preserves the spatial layout of the encoded image. As a consequence, by masking regions in the latent UV position map corresponding to regions we wish to modify, and by denoising the masked regions, one can apply intuitive edits to particular regions of the facial geometry. Please refer to RePaint (Lugmayr et al., 2022) for more details on the masking process. Even when using masks with sharp boundaries, the denoising process can take care of smoothly interpolating at the mask boundaries. This mask-based editing of facial geometries using our model is shown in Fig. 11. In the top row of Fig. 11 we mask the nose region of the latent position map, such that it remains fixed throughout the multiple steps of denoising. We then generate multiple geometry samples by varying the initial noise input to the diffusion model. The noise predicted at each denoising step is multiplied with the nose mask before being fed as input to the denoising UNet for the next time step. This denoising procedure leads to generations where the generated samples all share the same nose shape, but vastly differing facial identities. In the bottom row of Fig. 11, we show the result of inverse masking, where the face shape is held fixed while allowing the nose shape to change. Our model produces meaningful results in both cases.
The editing of facial geometry by masking the latent position map can also be guided further with user conditions. In Fig. 12, we show an example of an interactive sketching workflow, where an artist can progressively edit a generated geometry by modifying one region at a time. For this application, we start from an initial 3D geometry obtained from our model by conditioning it on a hand-drawn sketch (a). An artist further edits the sketch, such as to open the mouth for instance (b). We automatically mask the latent position map such that only the regions corresponding to the edited sketch lines are allowed to change. The masked regions of the latent position map are populated with noise and denoised through our model while using the edited sketch as the conditioning signal. When used in this fashion, our model can perform local edits on the initial geometry to reflect the changes made by the artist. Additional examples of how subsequently modifying the mouth (c), face shape (d) and nose (e) are translated to changes in the generated 3D face geometry are shown in the remaining columns of Fig. 12. Please see the supplemental video for another illustration of this workflow.
@@ 4.4. Dynamic Generation
Even though our model is only trained with static face shapes, we find that it can generate temporally stable 3D facial geometries when conditioned on per-frame FLAME parameters derived from animation sequences or on CLIP embeddings obtained from individual frames from in-the-wild videos. In Fig. 13, we show the generated 3D face geometry produced by our method when conditioned on various signals derived from videos. To demonstrate the use of sketches as dynamic conditioning, we use a recent face reconstruction technique (Chandran et al., 2023) to track the facial geometry in 3D from an in-the-wild video and then render out 2D sketches using a hand-painted texture map and the tracked geometry. We identify that the only pre-processing required to obtain temporally stable generations from CLIP embeddings is to smooth them with a box filter before using them as the conditioning signal. To further ensure stable generations across time, we use the same noise seed across the video as well as DDIM sampling. We kindly refer to our accompanying video for more dynamic results.
5. Conclusion
@@
We propose a new framework for 3D facial geometry generation based on latent diffusion models that can be guided using multiple types of conditionings. Our conditional geometry generator operates in a latent geometry space and can produce high resolution UV position maps. It can be seamlessly conditioned on hand-drawn sketches, 2D landmarks, Canny edges, FLAME-parameters, RGB portrait photos and even text; resulting in a comprehensive facial geometry generator that supports many applications like geometry super-resolution, geometry editing, face reconstruction, etc. We train our model from scratch on only static face shapes captured in a studio setting and yet demonstrate that our model can generalize reasonably to in-the-wild conditioning signals, and can also generate facial performances when conditioned on frames from video data. As limitations, we identify that our model can produce geometric artifacts and suboptimal geometry for extreme expressions, especially when controlled using FLAME’s jaw pose parameters. We are hopeful that our work will encourage future research towards building comprehensive facial foundation models that also consider facial geometry.

### v2

@@ 4 Results
We now show several results and applications of our new multimodal conditional 3D face geometry generation method. We begin by demonstrating control over the generated facial geometry using FLAME’s identity and expression parameters (Section 4.1). We then demonstrate multimodal conditioned geometry generation by guiding the denoising process using sketches, sparse 2D landmarks, Canny edges, portrait photos and text. We evaluate the effectiveness of these different modalities in guiding the generated geometry in Section 4.2. We also compare our method to the state-of-the-art related work both on text and image prompts in Section 4.2. In addition to using different conditioning modes, we show how one can also spatially restrict the guidance to a particular face region to perform precise geometry edits in Section 4.3. We can also generate dynamic facial performances by guiding our model from video inputs and demonstrate that our method can produce facial animations that are stable across time (Section 4.4). Finally, we discuss the limitations of our method in Section 4.5.
4.1 Identity and Expression Conditioning
@@
The base conditioning used to train our geometry generator are the identity and expression parameters from the FLAME model [3]. We use 300 identity parameters (), 100 expression parameters () and 3 jaw pose () parameters. We combine these FLAME parameters into a 403-dimensional conditioning vector which we pass through a 3-layer MLP with Leaky ReLU activation functions prior to injecting it into the diffusion model via cross-attention.
@@
Recollect that we do not use the geometry from the FLAME model itself to train our diffusion model. Instead we fit the FLAME model to the high quality facial geometry from Chandran et al. [14] only to obtain identity and expression parameters, and train the diffusion model directly on the geometry captured by Chandran et al.
@@
Disentangling Identity and Expression. Our diffusion model also preserves the disentanglement between facial identity and expression that is present in FLAME. In Fig. 4, we show how a smooth interpolation of FLAME’s identity and expression coefficients results in a smooth, yet nonlinear, interpolation of our generated geometry. We can observe that the facial expression remains fixed when interpolating between identities, and vice versa (please also refer to the supplemental video). To eliminate the randomness in the generation, we used the same initial noise to generate the interpolated geometries along with DDIM sampling [47].
4.2 Multimodal Conditioning
Beyond the underlying FLAME-based control, we introduce additional conditioning modes to control our diffusion model following Section 3.4. We now discuss the results of facial geometry generation by conditioning our diffusion model on sketches, sparse 2D landmarks, Canny edges, portrait RGB photos and text. Geometries generated for different conditionings from each of these additional modes are shown in Fig. 5.
@@ 4.2 Multimodal Conditioning
| 2D landmarks | 6.390 | 5.950 | 1.945 |
| Canny edges | 6.007 | 5.701 | 1.672 |
| Sketch | 5.521 | 5.106 | 1.792 |
| Portrait photo | 5.207 | 4.942 | 1.471 |
| FLAME parameters | 5.008 | 4.737 | 1.498 |
Quantitative Comparison. While our generated geometry follows the identity and expression seen in the input conditioning signal, the degree to which the conditioning signal constrains the generated geometry varies from mode to mode. We evaluate the effectiveness of each of our conditioning modes in guiding the generated geometry towards ground truth scans, by computing the Euclidean error between the generated geometry and the ground truth geometry for each conditioning mode. In Table 1, we report the vertex-to-vertex (V2V) error of 432 shapes from our validation set for each type of conditioning. We observe that conditioning signals that are more descriptive, such as FLAME parameters or portrait photos, obtain a lower error when compared to signals that are less descriptive of the final geometry (2D landmarks, Canny edges and sketches). Please refer to our supplementary material for error maps on our validation set.
Next, we compare our method to the state-of-the-art related work methods. First, we look at the text conditioning mode, because prompts from this mode are most commonly supported by related work. We compare our method to Describe3D [26], DreamFace [4] and two variants of FaceG2E [5], over 10 different text prompts. We report the average CLIP score (ViT-B/32) between the CLIP embeddings from the text prompt and the CLIP embeddings extracted from the rendered generated face geometry. Specifically, we evaluate text prompts that specify varied identities all with a neutral expression, because neutrals are supported by all methods. Each prompt is prepended with ”A shaded, textureless 3D face model of”, although for legibility, we shorten the text prompts when we add them to figures (e.g. Fig. 6 and Fig. 7). Please refer to our supplementary material for the exact text prompts. Among the compared methods, our method has the second highest CLIP score on text prompts that specify a neutral expression (Table 2).
Qualitative Comparison. We first show qualitative comparisons to state-of-the-art text-to-geometry methods on neutral text prompts in Fig. 6. Our method produces realistic faces that are subjectively on par with other techniques. Furthermore, it is important note that our method also natively supports generating faces with expressions by providing an expression description in the text prompt. This is a situation that other methods struggle with, as shown in Fig. 7.
Second, besides text prompts, DreamFace and our method also support RGB image prompts (Fig. 8). Our method achieves results comparable to DreamFace without relying on SDS optimization as part of the geometry generation. Note, that to aid the visual comparisons with previous methods, we complete the head in our results by deforming a template head to match our generated face. Despite training on purely studio data, we show how our model responds to conditionings derived from in-the-wild data in Fig. 9. For direct visual comparison with the identity and facial expression in the conditioning signal, we overlay the generated meshes onto validation images from whom the respective conditionings were extracted in Fig. 10.
Combining Two Conditionings. As our method involves learning additional modes of conditioning on top of an underlying FLAME conditioned diffusion model, we can also use more than one conditioning signal at inference time to guide the generation. In Fig. 11, we show how combining both FLAME parameter and portrait image conditioning lowers the vertex error on a validation sample, as the denoising UNet now has access to more information about the desired identity and expression. Note that only our base conditioning (FLAME parameters) is present when training any one of the other modalities cross-attention layers. Therefore, we can expect complementary results only when combining the FLAME parameter conditioning with another modality.
Inference time. We show that our method has significant inference-time benefits over its competitors that are mainly based on SDS optimization. It has the fastest average inference speed per sample on text-to-geometry generation, because once it is trained, it can directly sample 3D faces from the diffusion model. This setup alleviates the need for compute intensive iterative SDS optimization of a 3D face representation. Specifically, our method is more than seven times faster than its closest competitor FaceG2E (5.48 seconds vs. 42.87 seconds). To compare results with similar inference speeds, we run FaceG2E fast for approximately the same time as our own method and evaluate the CLIP score results (see Table 2).
@@
Controlling the Conditioning Strength. To control the strength of the guiding condition, we make use of classifier-free guidance [49] following Eq. 6. Increasing the guidance strength increases the effect that the input conditioning (prompt) has on the resulting geometry. Stronger guidance can lead to increased level of detail in the generated face geometry and greater resemblance with the input prompt. For example, in Fig. 12, the expression of the generated geometry of the subject in the first row displays stronger wrinkles, and a closer match to the portrait image when setting compared to setting it to . Unless specified differently, we use for all our conditional generation results.
@@
4.3 Geometry Editing
The latent space of our autoencoder preserves the spatial layout of the original UV position map, much like how the latent space of the image autoencoder in text-to-image models [9] preserves the spatial layout of the encoded image. As a consequence, by masking regions in the latent UV position map corresponding to regions we wish to modify, and by denoising the masked regions, one can apply intuitive edits to particular regions of the facial geometry. Please refer to RePaint [50] for more details on the masking process. Even when using masks with sharp boundaries, the denoising process can take care of smoothly interpolating at the mask boundaries. We show results of guiding the editing of facial geometry with user conditions in Fig. 13. Specifically, we show an interactive sketching workflow (6 seconds/sample), where an artist can progressively edit a generated geometry by modifying one region at a time.
@@ 4.4 Dynamic Generation
Although our model is only trained with static face shapes, we find that it can generate temporally stable 3D facial geometries when conditioned on per-frame FLAME parameters derived from animation sequences or on CLIP embeddings obtained from individual frames from in-the-wild videos. In Fig. 14, we show the generated 3D face geometry produced by our method when conditioned on various signals derived from videos. To demonstrate the use of sketches as dynamic conditioning, we use a recent face reconstruction technique [51] to track the facial geometry in 3D from an in-the-wild video and then render out 2D sketches using a hand-painted texture map. We identify that the only pre-processing required to obtain dynamically stable generations from CLIP embeddings is to temporally smooth them before using them as the conditioning signal. To further ensure stable generations across time, we use the same noise seed and DDIM sampling.
4.5 Limitations and Future Work
As limitations, we identify that our model can produce geometric artifacts for extreme expressions, especially when controlled using FLAME’s jaw pose parameters. This problem is mainly a data limitation and could be resolved by sourcing a larger dataset of extreme expressions, by oversampling expressions during training or by weighting the loss towards focusing more on extreme expressions. Additionally, we identify a limitation of extremely similar CLIP-based conditionings for left/right mirrors of asymmetrical face expressions, leading to a direction ambiguity in the geometry output. We refer to our supplementary material for further discussion and visualization of these failure cases. Beyond addressing those limitations, future work could incorporate facial appearance information into our method, enabling multimodal control over 3D faces with corresponding texture.
5 Conclusion
@@
We propose a new framework for 3D facial geometry generation based on a latent diffusion model that can be guided using multiple types of conditionings (prompts). Our conditional geometry generator operates in a latent geometry space. It can produce high quality geometry at comparably fast inference speeds using a UV position map representation. It can be seamlessly conditioned on hand-drawn sketches, 2D landmarks, Canny edges, FLAME-parameters, RGB portrait photos and text; resulting in a comprehensive facial geometry generator that supports many applications. For example, stochastic detail variation in the generated geometry or local geometry edits. We train our model from scratch on only static face shapes captured in a studio setting and yet demonstrate that our model can generalize reasonably to in-the-wild conditioning signals, and can also generate facial performances when conditioned on frames from video data.
Appendix A Ablation Studies
To validate the benefits of representing the face geometry as a delta from the template mesh, we train a latent diffusion model (LDM) on the full vertex map representation and a second LDM on the delta vertex map representation. We qualitatively compare both representations in Fig. 15 and can observe that the delta vertex map representation leads to less artifacts compared to the full vertex map representation (e.g.on the eyelids).
Next, we validate the benefits of adding geometry data augmentations to our training data. We compare FLAME parameter conditioned generations from two LDMs, where one was trained with and the other was trained without geometry data augmentations. As we illustrate in Table 3, our generations are closer to the ground truth geometry (lower vertex-to-vertex error), when using geometry data augmentations. This result indicates that using geometry data augmentations improves the models ability to capture unseen identities.
| V2V error | Mean | Median | Std |
| No augmentations | 4.093 | 3.692 | 2.170 |
| With augmentations | 3.757 | 3.352 | 2.085 |
Appendix B Implementation Details
For our dataset, we crop the full head face geometry to allocate more vertices to the face region, representing 50520 face vertices within each UV position map. At resolution we can represent reasonably high-resolution face geometry while being able to limit our VAE training time to 8 days using our training dataset of 7752 samples (1.4 seconds/iteration; batch size 8). We use a learning rate of 4.5e-6 and a codebook size of 8192. Note that for training our VAE, we did not use the data augmentations described in Section A as the autoencoder was already able to reconstruct test geometries with high accuracy when trained only on the studio dataset. The vertex error between reconstructions and the original geometry is usually below 0.3 millimeters and only very high-frequency details are lost. We visualize the VAE reconstruction error in Fig. 16. Next, we train our LDM for 4 days (1.6 seconds/iteration; batch size 12) with a learning rate of 1e-4 and diffusion timesteps . We utilize geometry data augmentations with corresponding FLAME fits during training (+200k samples) to allow for better generalization across identities during generation. Afterwards, we train each set of cross-attention layers with a learning rate of 1e-4 for 6 days (3.3 seconds/iteration; batch size 24) while keeping the LDM frozen. With a probability of 0.05, we randomly set either or or both to their null embeddings during training. This step enables classifier-free guidance at inference. Note that we do not add augmented geometry data to train the new cross-attention layers because we do not have access to paired mode-geometry data for modes such as portrait photos. We use the CLIP ViT-L/14 model [48] to extract 768 dimensional CLIP feature vectors as our conditioning representation for all modalities except the base FLAME parameter conditioning. We visualize the layers that pass our base FLAME parameter conditioning to the diffusion model in Fig. 17. These layers are trained jointly with the diffusion model parameters. Afterwards, the diffusion model and the base conditioning layers are frozen. For adding a new modality a newly added set of mode-specific layers are trained (linear layer, layer norm and a set of cross-attention layers). Sketches, portrait photos, Canny edges and landmarks are passed through a frozen CLIP image encoder before reaching their own mode-specific trainable layers. We train a different set of layers per mode (e.g. one for sketches, one for portrait photos etc.). Text is passed through a frozen CLIP text encoder before reaching text-specific trainable layers (Fig. 17). All training experiments were run on a single RTX A6000 GPU. Inference was run on single RTX A6000 GPUs, 3090 GPUs, and 1080 GPUs. Also note that the VAE and the LDM have to be trained only once and novel conditioning modes can be added by training only the new cross-attention layers. Unless mentioned otherwise, we generate every result by running DDPM sampling steps with conditioning strength . The average time to generate a geometry sample with our diffusion model is 6 seconds on a single 3090 GPU. To aid the visual similarity for the comparison with the state-of-the-art methods, we complete the head by deforming a template head to match our generated face. We run the CLIP score evaluation for all methods on full head renders. We align all 3D faces to the same space, before rendering each with the same camera. We use the CLIP ViT-B/32 variant for the score calculation and report the average score for each method.
Appendix C Exact Text Prompts
Table 4 lists the exact text prompts used for the CLIP score comparison and the figures in the main document.
| Nr. | Text prompt |
| 1 | A shaded, textureless 3D face model of an African woman with a neutral expression. |
| 2 | A shaded, textureless 3D face model of an overweight man with a neutral expression. |
| 3 | A shaded, textureless 3D face model of an old woman with a neutral expression. |
| 4 | A shaded, textureless 3D face model of a middle-aged Asian person with a neutral expression. |
| 5 | A shaded, textureless 3D face model of a middle-aged Caucasian woman with a neutral expression. |
| 6 | A shaded, textureless 3D face model of a woman with high cheekbones, a defined jawline, and a straight nose with a neutral expression. |
| 7 | A shaded, textureless 3D face model of a young woman with a round face, big eyes and small mouth with a neutral expression. |
| 8 | A shaded, textureless 3D face model of a man with a chubby face, a wide forehead and a wide nose with a neutral expression. |
| 9 | A shaded, textureless 3D face model of a young African man with a neutral expression. |
| 10 | A shaded, textureless 3D face model of a young Asian man with a neutral expression. |
| 11 | A shaded, textureless 3D face model of a smiling overweight man. |
| 12 | A shaded, textureless 3D face model of an overweight man shouting angrily. |
| 13 | A shaded, textureless 3D face model of a sad Caucasian man. |
| 14 | A shaded, textureless 3D face model of a middle-aged Asian person with a kiss face expression. |
| 15 | A shaded, textureless 3D face model of a smiling African woman. |
| 16 | A shaded, textureless 3D face model of a woman with high cheeckbones, a defined jawline, and a straight nose. Her mouth is opened to the side. |
| 17 | A shaded, textureless 3D face model of an angry Caucasian man. |
| 18 | A shaded, textureless 3D face model of a man with a chubby face, a wide forehead and a wide nose with a big smile on his face. |
| 19 | A shaded, textureless 3D face model of a young African man with a closed eyes facial expression. |
| 20 | A shaded, textureless 3D face model of a young Asian man with a very surprised facial expression. His eyes and mouth are wide open and the eyebrows raised. |
Appendix D Additional Geometry Editing Results
Further mask-based editing of facial geometries using our model is shown in Fig. 18. In the top row of Fig. 18 we mask the nose region of the latent position map, such that it remains fixed throughout the multiple steps of denoising. We then generate multiple geometry samples by varying the initial noise input to the diffusion model. The noise predicted at each denoising step is multiplied with the nose mask before being fed as input to the denoising UNet for the next time step. This denoising procedure leads to generations where the generated samples all share the same nose shape, but vastly differing facial identities. In the bottom row of Fig. 18, we show the result of inverse masking, where the face shape is held fixed while allowing the nose shape to change. Our model produces meaningful results in both cases.
Appendix E Additional Quantitative and Qualitative Results
For the base FLAME parameter conditioning, we visualize the error maps to the ground truth scanned geometry in Fig. 19. Next, we compare the text-to-geometry generation results of HeadArtist [6] and HumanNorm [20] with our method. Both are based on Deep Marching Tetrahedra [21] and SDS optimization [22] and can represent face parts beyond the skin. The extracted face geometries differ in topology and optimizing for one sample takes around one hour on a 3090 GPU. In contrast, our method’s inference speed is 1000-times faster on a 3090 GPU and produces results in a single common topology.
Appendix F Failure Cases
We do observe geometric artifacts around the mouth region for extreme expressions (Figure 20, column 2), due to limited extreme expressions in our training data. Additionally, the generated face geometry can open the mouth to the wrong side when conditioning with CLIP embeddings (Figure 20, column 3 and 4). We identify that this behavior occurs, when the CLIP embeddings for both mouth opening directions (left/right) are extremely similar (cosine similarity close to 1). Thus, this behavior is caused by the similarity of specific conditionings and not by the diffusion model.